OpenStudy (anonymous):

graph the equation by plotting points y=1/4x-4. complete the ordered pairs. (4,__), (0,__)

OpenStudy (sandra):

ok, so you can graph it also

OpenStudy (sandra):

but what they're giving you in this problem is the x coordinate

OpenStudy (sandra):

so to complete the coordinate, you insert each x value from those coordinates into your equation

OpenStudy (sandra):

so for the two coordinates (x,y): 1. (4,_) , x = 4, and y= 1/4*4 - 4 = -3. so the coordinate is (4,-3) 2. (0,_), x= 0, and y= 1/4*0 - 4 = -4. so the coordinate is (0,-4)